# A replica analysis of under-bagging

### v1

@@ 3 Asymptotics of the estimator for the randomized cost function
In this section, we present the asymptotic properties of the estimators (4) and (5) for the randomized cost functions. It is described by a small finite set of scalar quantities determined as a solution of nonlinear equations, which we refer to as self-consistent equations.
First, let us define the self-consistent equations that determines a set of scalar quantities $B\in\mathbb{R},\Theta=(q,m,\chi,v)\in(0,\infty)\times\mathbb{R}\times(0,\infty)\times(0,\infty)$ and $\hat{\Theta}=(\hat{Q},\hat{m},\hat{\chi},\hat{v})\in(0,\infty)\times\mathbb{R}\times(0,\infty)\times(0,\infty)$ that characterize the statistical behavior of (4) and (5).
Claim 1 (Self-consistent equations)
The set of quantities $\Theta$ and $\hat{\Theta}$ are given as the solution of the following set of non-linear equations that are referred to as self-consistent equations. Let $\hat{\mathsf{w}}$ and $\hat{\mathsf{u}}^{\pm}$ be the solution of the following one-dimensional randomized optimization problems:
@@
That is, $q,m,v,B$ represents the squared norm of the averaged weight vector, the inner product with the cluster center, the variance with respect to the resampling, and the bias, respectively.
@@
Let $(B,m,q,v)$ be the solution of the self-consistent equation in Claim 1. Also, let $l=0,1,\dots\in{\mathbb{N}}$ be the non-negative integer. Also, $\hat{u}^{\pm}(\mathbf{\boldsymbol{c}},D)\equiv\mathbf{\boldsymbol{x}}^{\pm}\cdot\hat{\mathbf{\boldsymbol{w}}}(\mathbf{\boldsymbol{c}},D)/\sqrt{N}+\hat{B}(\mathbf{\boldsymbol{c}},D)$ and $\hat{u}^{\pm}(\mathbf{\boldsymbol{c}},D)\equiv\mathbf{\boldsymbol{x}}^{\pm}\cdot\hat{\mathbf{\boldsymbol{w}}}(\mathbf{\boldsymbol{c}},D)/\sqrt{N}+\bar{B}$ be the predictions of the estimators (4) and (5) for an input $\mathbf{\boldsymbol{x}}^{\pm}$ which is generated as
@@
$\mathbb{E}_{D,\mathbf{\boldsymbol{x}}^{\pm}}\left[\mathbb{E}_{\mathbf{\boldsymbol{c}}}[\psi(\hat{u}^{\pm})]^{l}\right]=\mathbb{E}_{\xi_{u}\sim{\mathcal{N}}(0,\Delta)}\left[\mathbb{E}_{\eta_{u}\sim{\mathcal{N}}(0,\Delta)}[\psi(B\pm m+\sqrt{q}\xi_{u}+\sqrt{v}\eta_{u})]^{l}\right],$ (31)
@@
This result indicates that the fluctuation with respect to resampling conditioned by $(\mathbf{\boldsymbol{x}},y)$ is effectively described by the Gaussian random variable $\eta_{u}$, and the fluctuation with respect to the noise term in $\mathbf{\boldsymbol{x}}$ is described by the Gaussian random variable $\xi_{u}$. Schematically, this behavior can be represented as follows:
$\hat{u}^{\pm}=\underset{\rm bias}{\uwave{B}}+\underset{\begin{subarray}{c}\rm correlation~{}with\\ \rm the~{}cluster~{}center\end{subarray}}{\uwave{(\pm m)}}+\underset{\begin{subarray}{c}{\rm fluctuation~{}from}\\ {\rm the~{}noise~{}}\mathbf{\boldsymbol{z}}\end{subarray}}{\uwave{\sqrt{q}\xi_{u}}}+\underset{\begin{subarray}{c}{\rm fluctuation~{}from}\\ {\rm random~{}reweighting~{}}\mathbf{\boldsymbol{c}}\\ {\rm conditioned~{}by~{}}\mathbf{\boldsymbol{x}}\end{subarray}}{\uwave{\sqrt{v}\eta_{u}}}.$ (32)
@@
Let $(B,m,q,v)$ be the solution of the self-consistent equation in Claim 1. Then, the distribution of the prediction $\hat{u}_{K}^{\pm}$ averaged over $K$-realization of the reweighing coefficient $\{\mathbf{\boldsymbol{c}}_{k}\}_{k=1}^{K}$, when the noise term $\mathbf{\boldsymbol{z}}$ fluctuates, is equal to that of the following random quantity:
$\hat{u}^{\pm}_{K}=B\pm m+\sqrt{q}\xi_{u}+\sqrt{\frac{v}{K}}\eta_{u},\quad\xi_{u},\eta_{u}\sim{\mathcal{N}}(0,\Delta).$ (33)
@@
Figure 1: Comparison between the empirical distribution of $\hat{u}_{K}^{\pm}$ (histogram), which is obtained by a single realization of the training data $D$ of finite size, and the theoretical prediction in Claim 4 (dashed). Different colors represent resampling averages with different numbers of realizations of reweighting coefficient $\mathbf{\boldsymbol{c}}$.
@@
Let $(m,q,B)$ be the solution of the self-consistent equations in Claim 1. Then, for the prediction $\hat{u}_{K}^{\pm}$ averaged over $K$-realization of the reweighing coefficient $\{\mathbf{\boldsymbol{c}}_{k}\}_{k=1}^{K}$, specificity ${\mathcal{S}}$, recall ${\mathcal{R}}$, and the $F$-measure ${\mathcal{F}}$ that is obtained as the weighted harmonic mean of specificity and recall, is given as follows:
@@
If the bias is properly estimated, the generalization performance is determined by the variance of the prediction $\Delta(q+v/K)$.
@@ 3.1 Cross-checking with numerical experiments
To check the validity of Claim 4, which is the most important result for characterizing the generalization performance, we briefly compare the result of numerical solution of the self-consistent equations with numerical experiments of finite-size systems.
@@
We consider the case of subsampling (6)-(7). For the loss function, we use the cross entropy loss; $l^{+}(x)=-\log(\sigma(x))$ and $l^{-}(x)=-\log(1-\sigma(x))$, where $\sigma(\cdot)$ is the sigmoid function. The regularizer is the ridge: $r(x)=x^{2}/2$. Also, the bias is estimated by minimizing the loss as in (4). In experiments, the scikit-learn package is used (Pedregosa et al., 2011), and the system sizes $N=2^{10}$ and $N=2^{13}$ are used. For simplicity, we fix the regularization parameter as $\lambda=10^{-4}$, the size of the training data as $(M^{+}+M^{-})/N=1/2$, and the variance of the noise as $\Delta=0.75^{2}$. For evaluating the distribution of $\hat{u}_{K}^{\pm}$ empirically, we use $10^{5}$ examples of $\mathbf{\boldsymbol{x}}^{\pm}$.
@@
Figure 1 shows the comparison the distribution in Claim 4 and the empirical distribution obtained by experiments with finite $N$. Different colors corresponds to different numbers of samples for the resampling average. Each panel corresponds to a different set of $(M^{+}/M,y,N)$, where $y$ is the label to indicate positive ($y=1$) or negative ($y=-1$) classes. In all cases, although some discrepancies between theoretical prediction and experiment are observed when $N=2^{10}$, they are in agreement when $N=2^{13}$, demonstrating the validity of Claim 4.
@@ 4 Under-bagging and regularization
Figure 2: The comparison of the dependence of the $F$-measure ${\mathcal{F}}$ on $\alpha^{-}-\alpha^{+}$, which is the difference in the size between the majority (negative) and the minority (positive) classes, between US (dashed) and UB (solid). Each panel corresponds to different values of $(\Delta,\lambda)$. Different colors of the lines represent the result for different sizes of the minority classes $\alpha^{+}$.
In this section, using Claim 4, we compare the performance of several estimators, that is, (i) US, (ii) UB, and (iii) SW.
For simplicity, we focus on the case of ridge-regularized cross-entropy loss as in subsection 3.1. When training a model using an under-sampled data set as in (6)-(7), we can expect that the estimated bias term $\hat{B}(\mathbf{\boldsymbol{c}},D)$ to be zero, since the number of positive and negative samples is balanced. Therefore, we set the bias as zero a priori when using subsampled data sets. Furthermore, as investigated in the previous studies (Dobriban & Wager, 2018; Mignacco et al., 2020), when training a linear classifier using balanced data and a ridge-regularized cross-entropy loss, an infinitely large large regularization parameter yields the best direction for the classification plane. However, as reported in (Mignacco et al., 2020; Takahashi, 2024), the discrepancy between the theory and experiment with finite N can be huge if too a large regularization parameter is used. Therefore, for US and UB, we consider several finite regularization parameters for US and UB. When considering the weighting, the behavior of bias and the optimal regularization parameter, the bias term is estimated as in (4) and the regularization parameter is optimized to maximize the $F$-measure using the Nelder-Mead method in Optim.jl package (Mogensen & Riseth, 2018).

### v2

@@ 3 Asymptotics of the estimator for the randomized cost function
In this section, we present the asymptotic properties of the estimators (4) and (5) for the randomized cost functions. It is described by a small finite set of scalar quantities determined as a solution of nonlinear equations, which we refer to as self-consistent equations. The derivation is outlined in Appendix A.
As noted in the related works section, the salient feature of the asymptotics in the LSL is that the macroscopic properties of the learning results, such as the predictive distribution, typically do not depend on each realization of the training data. In this paper, we assume these concentration properties in advance to derive the formulas for the asymptotic properties.
Assumption 1 (Concentration of the macroscopic properties of the learning results)
We assume that the macroscopic properties of the estimators (4)-(5) will concentrate at LSL. Concretely, we assume that the values of $(N^{-1}\sum_{i}\mathbb{E}_{\mathbf{\boldsymbol{c}}}[\hat{w}_{i}(\mathbf{\boldsymbol{c}},D)]^{2},N^{-1}\sum_{i}\mathbb{E}_{\mathbf{\boldsymbol{c}}}[\hat{w}_{i}(\mathbf{\boldsymbol{c}},D)],N^{-1}\sum_{i}{\mathbb{V}}_{\mathbf{\boldsymbol{c}}}[\hat{w}_{i}(\mathbf{\boldsymbol{c}},D)],\hat{B}(\mathbf{\boldsymbol{c}},D))$ does not fluctuate with respect to the realization of $D$ at LSL. Also, Let $\mathbf{\boldsymbol{x}}^{\pm}=\pm\mathbf{\boldsymbol{v}}/\sqrt{N}+\mathbf{\boldsymbol{z}},~{}z_{i}\sim_{\rm iid}p_{z},i\in[N]$ be a new input. We assume that the distribution of the prediction for this new input $\hat{s}^{\pm}(\mathbf{\boldsymbol{c}},D)=\mathbf{\boldsymbol{x}}^{\pm}\cdot\hat{\mathbf{\boldsymbol{w}}}(\mathbf{\boldsymbol{c}},D)/\sqrt{N}+\hat{B}(\mathbf{\boldsymbol{c}},D)$ (or $\bar{B}$ when the bias is fixed) with respect to $\mathbf{\boldsymbol{c}}$ and $\mathbf{\boldsymbol{z}}$ does not fluctuate with respect to the realization of $D$ at LSL.
Under the above assumption, we obtain the following prediction of the behavior of the estimators (4)-(5).
First, let us define the self-consistent equations that determines a set of scalar quantities $B\in\mathbb{R},\Theta=(q,m,\chi,v)\in(0,\infty)\times\mathbb{R}\times(0,\infty)\times(0,\infty)$ and $\hat{\Theta}=(\hat{Q},\hat{m},\hat{\chi},\hat{v})\in(0,\infty)\times\mathbb{R}\times(0,\infty)\times(0,\infty)$ that characterize the statistical behavior of (4) and (5). The readers interested in the implications of these quantities can skip Definition 1 and Claim 1, and proceed to Claim 2.
Definition 1 (Self-consistent equations)
The set of quantities $\Theta$ and $\hat{\Theta}$ are defined as the solution of the following set of non-linear equations that are referred to as self-consistent equations. Let $\hat{\mathsf{w}}$ and $\hat{\mathsf{u}}^{\pm}$ be the solution of the following one-dimensional randomized optimization problems:
@@
That is, $q,m,v,B$ represents the squared norm of the averaged weight vector, the inner product with the cluster center, the variance with respect to the resampling, and the bias, respectively, at LSL.
@@
Let $(B,m,q,v)$ be the solution of the self-consistent equation in Definition 1, and let $l=1,2,\dots\in{\mathbb{N}}$ be the non-negative integer. Also, $\hat{s}^{\pm}(\mathbf{\boldsymbol{c}},D)\equiv\mathbf{\boldsymbol{x}}^{\pm}\cdot\hat{\mathbf{\boldsymbol{w}}}(\mathbf{\boldsymbol{c}},D)/\sqrt{N}+\hat{B}(\mathbf{\boldsymbol{c}},D)$ and $\hat{s}^{\pm}(\mathbf{\boldsymbol{c}},D)\equiv\mathbf{\boldsymbol{x}}^{\pm}\cdot\hat{\mathbf{\boldsymbol{w}}}(\mathbf{\boldsymbol{c}},D)/\sqrt{N}+\bar{B}$ be the predictions of the estimators (4) and (5) for an input $\mathbf{\boldsymbol{x}}^{\pm}$ which is generated as
@@
$\mathbb{E}_{D,\mathbf{\boldsymbol{x}}^{\pm}}\left[\mathbb{E}_{\mathbf{\boldsymbol{c}}}[\psi(\hat{s}^{\pm})]^{l}\right]=\mathbb{E}_{\xi_{u}\sim{\mathcal{N}}(0,\Delta)}\left[\mathbb{E}_{\eta_{u}\sim{\mathcal{N}}(0,\Delta)}[\psi(B\pm m+\sqrt{q}\xi_{u}+\sqrt{v}\eta_{u})]^{l}\right],$ (31)
@@
This result indicates that the fluctuation with respect to resampling conditioned by $\mathbf{\boldsymbol{x}}$ is effectively described by the Gaussian random variable $\eta_{u}$, and the fluctuation with respect to the noise term in $\mathbf{\boldsymbol{x}}$ is described by the Gaussian random variable $\xi_{u}$. Schematically, this behavior can be represented as follows:
$\hat{s}^{\pm}=\underset{\rm bias}{\uwave{B}}+\underset{\begin{subarray}{c}\rm correlation~{}with\\ \rm the~{}cluster~{}center\end{subarray}}{\uwave{(\pm m)}}+\underset{\begin{subarray}{c}{\rm fluctuation~{}from}\\ {\rm the~{}noise~{}}\mathbf{\boldsymbol{z}}\end{subarray}}{\uwave{\sqrt{q}\xi_{u}}}+\underset{\begin{subarray}{c}{\rm fluctuation~{}from}\\ {\rm random~{}reweighting~{}}\mathbf{\boldsymbol{c}}\\ {\rm conditioned~{}by~{}}\mathbf{\boldsymbol{x}}\end{subarray}}{\uwave{\sqrt{v}\eta_{u}}}.$ (32)
@@
Let $(B,m,q,v)$ be the solution of the self-consistent equation in Definition 1. Then, the distribution of $\hat{s}_{K}^{\pm}=\frac{1}{K}\sum_{k=1}^{K}\mathbf{\boldsymbol{x}}\cdot\hat{\mathbf{\boldsymbol{w}}}(\mathbf{\boldsymbol{c}}_{k},D)/\sqrt{N}+\hat{B}(\mathbf{\boldsymbol{c}}_{k},D)$ (or $\bar{B}$ when the bias is fixed), the prediction averaged over $K$-realization of the reweighing coefficient $\{\mathbf{\boldsymbol{c}}_{k}\}_{k=1}^{K}$ when the noise term $\mathbf{\boldsymbol{z}}$ fluctuates, is equal to that of the following random quantity:
$\hat{s}^{\pm}_{K}=B\pm m+\sqrt{q}\xi_{u}+\sqrt{\frac{v}{K}}\eta_{u},\quad\xi_{u},\eta_{u}\sim{\mathcal{N}}(0,\Delta).$ (33)
@@
Figure 1: Comparison between the empirical distribution of $\hat{s}_{K}^{\pm}$ (histogram), which is obtained by a single realization of the training data $D$ of finite size, and the theoretical prediction in Claim 3 (dashed). Different colors represent resampling averages with different numbers of realizations of reweighting coefficient $\mathbf{\boldsymbol{c}}$.
The US, which uses a single realization of the resampled data set, corresponds to the case of $K=1$.
@@
Let $(q,m,v,B)$ be the solution of the self-consistent equations in Definition 1. Then, for $\hat{s}_{K}^{\pm}=\frac{1}{K}\sum_{k=1}^{K}\mathbf{\boldsymbol{x}}\cdot\hat{\mathbf{\boldsymbol{w}}}(\mathbf{\boldsymbol{c}}_{k},D)/\sqrt{N}+\hat{B}(\mathbf{\boldsymbol{c}}_{k},D)$ (or $\bar{B}$ when the bias is fixed), the prediction averaged over $K$-realization of the reweighing coefficient $\{\mathbf{\boldsymbol{c}}_{k}\}_{k=1}^{K}$, the specificity ${\mathcal{S}}$, the recall ${\mathcal{R}}$, and the $F$-measure ${\mathcal{F}}$ are given as follows:
@@
We remark that for any $K\geq 1$, the variable $m$, which represents the correlation between the cluster center $\mathbf{\boldsymbol{v}}$ and the weight vector $\hat{\mathbf{\boldsymbol{w}}}(\mathbf{\boldsymbol{c}},D)$ is a constant, that is, the bagging procedure just reduces the variance of the prediction and does not improves the direction of the classification plane. An important consequence is that for any regularization parameter $\lambda>0$, UB with $K>1$ is a higher value of the $F$-measure when $B=0$ and $m>0$, indicating the superiority of UB over US.
@@ 3.1 Cross-checking with numerical experiments
To check the validity of Claim 3, which is the most important result for characterizing the generalization performance, we briefly compare the result of numerical solution of the self-consistent equations with numerical experiments of finite-size systems. We also compare the values of $(q,m,v,B)$ with $(N^{-1}\sum_{i}\mathbb{E}_{\mathbf{\boldsymbol{c}}}[\hat{w}_{i}(\mathbf{\boldsymbol{c}},D)]^{2},N^{-1}\sum_{i}\mathbb{E}_{\mathbf{\boldsymbol{c}}}[\hat{w}_{i}(\mathbf{\boldsymbol{c}},D)],N^{-1}\sum_{i}{\mathbb{V}}_{\mathbf{\boldsymbol{c}}}[\hat{w}_{i}(\mathbf{\boldsymbol{c}},D)],\hat{B}(\mathbf{\boldsymbol{c}},D))$ obtained by numerical experiments with finite $N$, that is, the right-hand-side of equations in Claim 1 but with finite $N$.
@@
We consider the case of subsampling (6)-(7). For the loss function, we use the cross-entropy loss; $l^{+}(x)=-\log(\sigma(x))$ and $l^{-}(x)=-\log(1-\sigma(x))$, where $\sigma(\cdot)$ is the sigmoid function. The regularizer is the ridge: $r(x)=x^{2}/2$. Also, the bias is estimated by minimizing the loss as in (4). In experiments, the scikit-learn package is used (Pedregosa et al., 2011). For simplicity, we fix the size of the overall training data as $(M^{+}+M^{-})/N=1/2$. For evaluating the distribution of $\hat{s}_{K}^{\pm}$ empirically, we use $10^{5}$ examples of $\mathbf{\boldsymbol{x}}^{\pm}$.
Predictive distributions
@@
Figure 1 shows the comparison between the distribution in Claim 3 and the empirical distribution obtained by experiments with single realization of the training data $D$ of finite size $N$. Different colors corresponds to different numbers of samples for the resampling average. Each panel corresponds to a different set of $(M^{+}/M,y,N)$, where $y$ is the label to indicate positive ($y=1$) or negative ($y=-1$) classes. For simplicity, we fix the regularization parameter as $\lambda=10^{-4}$, and the variance of the noise as $\Delta=0.75^{2}$. For numerical experiments, the system sizes $N=2^{10}$ and $N=2^{13}$ are used. In all cases, they are in agreement when $N=2^{13}$, demonstrating the validity of Claim 3 and Assumption 1, although some discrepancies between theoretical prediction and experiment are observed when $N=2^{10}$.
(a) $(\sqrt{\Delta},\lambda)=(1.5,10^{-5})$
(b) $(\sqrt{\Delta},\lambda)=(1.5,10^{-3})$
(c) $(\sqrt{\Delta},\lambda)=(1.5,10^{-1})$
(d) $(\sqrt{\Delta},\lambda)=(0.75,10^{-5})$
(e) $(\sqrt{\Delta},\lambda)=(0.75,10^{-3})$
(f) $(\sqrt{\Delta},\lambda)=(0.75,10^{-1})$
(g) $(\sqrt{\Delta},\lambda)=(0.5,10^{-5})$
(h) $(\sqrt{\Delta},\lambda)=(0.5,10^{-3})$
(i) $(\sqrt{\Delta},\lambda)=(0.5,10^{-1})$
Figure 2: Comparison of the comparison of $(q,m,v,B)$, obtained as the solution of the self-consistent equation (lines), and $(N^{-1}\sum_{i}\mathbb{E}_{\mathbf{\boldsymbol{c}}}[\hat{w}_{i}(\mathbf{\boldsymbol{c}},D)]^{2},N^{-1}\sum_{i}\mathbb{E}_{\mathbf{\boldsymbol{c}}}[\hat{w}_{i}(\mathbf{\boldsymbol{c}},D)],N^{-1}\sum_{i}{\mathbb{V}}_{\mathbf{\boldsymbol{c}}}[\hat{w}_{i}(\mathbf{\boldsymbol{c}},D)],\hat{B}(\mathbf{\boldsymbol{c}},D))$ (markers with error bars). Reported experimental results are averaged over several realization of $D$ depending on the size $N$. The error bars represent standard deviations. Each panel corresponds to different values of $(\Delta,\lambda)$. Different colors of the lines represent the result for different system sizes $N$.
Parameters $\Theta$
Figure 2 shows the comparison of $(q,m,v,B)$, obtained as the solution of the self-consistent equation, and $(N^{-1}\sum_{i}\mathbb{E}_{\mathbf{\boldsymbol{c}}}[\hat{w}_{i}(\mathbf{\boldsymbol{c}},D)]^{2},N^{-1}\sum_{i}\mathbb{E}_{\mathbf{\boldsymbol{c}}}[\hat{w}_{i}(\mathbf{\boldsymbol{c}},D)],N^{-1}\sum_{i}{\mathbb{V}}_{\mathbf{\boldsymbol{c}}}[\hat{w}_{i}(\mathbf{\boldsymbol{c}},D)],\hat{B}(\mathbf{\boldsymbol{c}},D))$. The values are plotted as functions of $\alpha^{-}-\alpha^{+}$, the difference of the normalized size of the majority and the minority classes. Each panel corresponds to a different set of $(\lambda,\sqrt{\Delta})$. For numerical experiments, two different system sizes $N=256(=2^{8})$, and $2048(=2^{11})$ are used. To take the average over $\mathbf{\boldsymbol{c}}$, $128$ independent realizations are used. Reported experimental results are averaged over several realization of $D$ depending on the size $N$. The results show that the theoretical predictions and the experimental values are in good agreement. Also, the error bars represent standard deviations that decrease as the system size grows, indicating the concentration of these quantities. Overall, the figure confirms the validity of Claim 1 and Assumption 1.
@@ 4 Under-bagging and regularization
Figure 3: Comparison of the dependence of the $F$-measure ${\mathcal{F}}$ on $\alpha^{-}-\alpha^{+}$, which is the difference in the size between the majority (negative) and the minority (positive) classes, between US (dashed) and UB (solid). Each panel corresponds to different values of $(\Delta,\lambda)$. Different colors of the lines represent the result for different sizes of the minority classes $\alpha^{+}$.
In this section, using Claim 4, we compare the performance of several estimators, that is, US, UB, and SW.
For simplicity, we focus on the case of ridge-regularized cross-entropy loss as in subsection 3.1. When training a model using an under-sampled data set as in (6)-(7), we can expect that the estimated bias term $\hat{B}(\mathbf{\boldsymbol{c}},D)$ to be zero, since the number of positive and negative samples is balanced. See also Figure 2. Therefore, we set the bias to zero a priori when using subsampled data sets. Furthermore, as investigated in the previous studies (Dobriban & Wager, 2018; Mignacco et al., 2020), when training a linear classifier using balanced data and a ridge-regularized cross-entropy loss, an infinitely large regularization parameter yields the best direction for the classification plane. However, as reported in (Mignacco et al., 2020; Takahashi, 2024), the discrepancy between the theory and experiment with finite $N$ can be huge if too a large regularization parameter is used because the norm of the weight vector can be pathologically small. Therefore, for US and UB, we consider several finite and fixed regularization parameters. When considering the weighting, the behavior of the bias and the optimal regularization parameter is not known a priori. Hence the bias term is estimated as in (4) and the regularization parameter is optimized to maximize the $F$-measure using the Nelder-Mead method in Optim.jl package (Mogensen & Riseth, 2018).